### v2

@@ Abstract
Text-to-video diffusion models have shown remarkable progress in generating coherent video clips from textual descriptions. However, the interplay between motion, structure, and identity representations in these models remains under-explored. Here, we investigate how self-attention query features (a.k.a. $Q$ features) simultaneously govern motion, structure, and identity and examine the challenges arising when these representations interact. Our analysis reveals that Q affects not only layout, but that during denoising Q also has a strong effect on subject identity, making it hard to transfer motion without the side-effect of transferring identity. Understanding this dual role enabled us to control query feature injection (Q injection) and demonstrate two applications: (1) a zero-shot motion transfer method that is 20$\times$ more efficient than existing approaches, and (2) a training-free technique for consistent multi-shot video generation, where characters maintain identity across multiple video shots while Q injection enhances motion fidelity.
Project page: https://research.nvidia.com/lab/par
/MotionByQueries/
@@
1 Introduction
Video generation from text is at the forefront of generative AI. Despite great progress in controlling entities in video, several major challenges remain, including generating natural and engaging motion and preserving consistent identity of entities throughout the video. These two goals tend to form a trade-off: It is easy to preserve consistency if motion is strongly limited, and making entities move makes it harder to enforce consistency, because the appearance of an entity changes. A major challenge lies in understanding how motion and identity are represented in various components of video generation models and how to effectively control them.
This limited understanding hinders downstream applications in video generation. For instance, current motion transfer approaches Yatim et al. (2024); Zhao et al. (2024); Wang et al. (2024a) still rely on tuning or test-time optimization, while their text-to-image layout transfer counterparts already operate through inference-time feature manipulation Cao et al. (2023); Alaluf et al. (2023). Better model understanding could potentially lead to similar strides here. As another example, consider consistent characters in multi-shot video generation, where the goal is to preserve consistency of character identity and appearance across various shots. Image-based models tackle this task through feature-sharing, but applying the same ideas to video leads to loss of motion because the shared features encode both identity and motion.
To understand representation of motion and identity in a video model, let us first draw analogies from text-to-image (T2I) models which are better understood. Motion can be viewed as a ”3D-shape” in the 3D tensor defined by a sequence of frames, so it is natural to look at representation of shape and structure in image generation models.
Previous works showed that diffusion-based T2I models establish the layout of the generated image in early generation steps (Patashnik et al., 2023). Several studies Cao et al. (2023); Alaluf et al. (2023); Tewel et al. (2024) have analyzed how different model components determine this structure. They showed that self-attention queries (Q) encode structural layout information. They found that injecting queries between images during generation “copies” shape while preserving appearance. As motion is the video equivalent of structure, we investigate the relationship between query vectors, motion and identity in video generation.
@@ 1 Introduction
For motion transfer, we find that injecting Q features from a source video during denoising of a generated video, allows to transfer the motion to a new video in a zero-shot manner (without any additional fine-tuning or optimization). Our simple pipeline achieves generation quality close to leading methods, while being 20$\times$ more efficient than existing approaches.
@@
Our main contributions: 1) We provide the first systematic analysis of Q-features in text-to-video diffusion models, revealing their dual role in encoding both motion and identity information, with effects persisting longer into the denoising process. 2) We introduce “Motion by Queries”, a zero-shot motion transfer approach that achieves comparable quality to current methods while being 20$\times$ more efficient. 3) We present the first training-free method for consistent multi-shot video generation that balances character consistency and motion quality.
Figure 2: Same-class motion transfer suffers from identity leakage (purple), which worsens with increased Q injection. Cross-class transfer (green) achieves reasonable separation at 40% injection, where motion quality is also preserved. The leftmost purple data point corresponds to results for random images of the same-class, as no injection occurs. Quantitative results were obtained using the motion-transfer data of Wang et al. (2024a). For illustration we show frames of the videos in Fig. 1.
@@ 2 Related work
Text-to-video synthesis. Following large, diffusion-based (Ho et al., 2020) text-to-image models (Rombach et al., 2021; Ramesh et al., 2022), works sought to replicate their success in video generation. Early text-based video approaches (Ho et al., 2022) leveraged cascaded approaches for time and space super-resolution. Methods commonly leveraged pre-trained text-to-image models’ knowledge, expanding them into video models (Wang et al., 2023a; He et al., 2022; Blattmann et al., 2023; Zhou et al., 2022; Wang et al., 2023c; Singer et al., 2023; Luo et al., 2023; Ge et al., 2023; Zhang et al., 2023a; Bar-Tal et al., 2024). Concurrent to such approaches, image-to-video models emerged as powerful alternatives (Gu et al., 2023; Wang et al., 2023b; Zhang et al., 2023b). While not strictly text-conditioned, these can be paired with a text-to-image model to generate an initial frame, which is then animated. Our work builds on existing T2V models (Chen et al., 2024), enabling them to maintain consistent characters across individually generated scenes.
@@
Motion Transfer in Video Generation. Motion transfer aims to guide video generation using motion patterns extracted from a source video, spanning camera movements and object deformations while maintaining structural integrity. Recent approaches commonly require model adaptation, either through fine-tuning temporal attention layers (Jeong et al., 2024; Ren et al., 2024), employing dual-path LoRA architectures to separate appearance and motion learning (Zhao et al., 2024), or learning specialized motion embeddings (Wang et al., 2024a). Alternative methods utilize test-time optimization to guide the generation process (Yatim et al., 2024). In contrast to these approaches that require optimization or fine-tuning steps, our method enables zero-shot motion transfer through query injection, requiring no additional training or optimization.
Video Editing. While motion transfer focuses on replicating specific motion patterns from source to target videos, video editing methods focus on modifying specific attributes of an input video while preserving its other characteristics. Recent approaches employ various mechanisms for content modification: Bai et al. (2024) is also based on feature injection. However, it focuses on very specific attribute changes while preserving appearance and lacks the broader context of our Q injection analysis. Zhao et al. (2023) combine ControlNet with CLIP-based guidance. Earlier works adapted text-to-image models for video editing through different attention manipulation strategies, such as injecting complete self-attention maps (Qi et al., 2023). While this direction spawned numerous works (Liu et al., 2024; Shin et al., 2024; Cong et al., 2023; Wu et al., 2023), these image-based approaches cannot faithfully capture motion patterns without replicating the source appearance, as the underlying models lack the capacity to process temporal dynamics independently.
Consistent generation aims to maintain consistent subjects across outputs produced by a generative model. This task has typically been considered under through the lens of text-to-image generation. A common approach is to leverage personalization (Gal et al., 2022; Ruiz et al., 2022) to promote consistency, either through inptaining with a personalized model (Jeong et al., 2023), by iteratively generating multi-character images using personalized LoRA models (Ryu, 2023), or by clustering randomly generated images and training LoRAs for large, semi-consistent clusters (Avrahami et al., 2023). Rather than fine-tuning a model, an encoder (Ye et al., 2023; Wei et al., 2023; Gal et al., 2023) can be used to inject an identity at inference time, but encoders require pre-training on large datasets, and struggle to accurately generalize to arbitrary domains. Similar issues arise when working with models that tune the model on storyboard datasets in order to augment it with additional conditioning on sets of image frames (Feng et al., 2023; Liu et al., 2023). Most recently, works (Tewel et al., 2024; Fan et al., 2024) explored character consistency without personalization, employing feature sharing approaches to generate consistent characters across image batches, without tuning or pre-training.
Structure and Appearance in Text-to-Image Models Early work in text-to-image generation revealed distinct roles for different components of attention mechanisms in diffusion-based models. Tewel et al. (2023) demonstrated that attention maps, derived from Keys and Queries, act as a “Where” pathway controlling the compositional structure, while Values serve as a “What” pathway determining visual appearance. Further studies established that the layout of generated images is primarily determined in the initial denoising steps (Patashnik et al., 2023). Through detailed analysis of model components,  (Alaluf et al., 2023; Tewel et al., 2024) identified self-attention queries as crucial encoders of structural layout information. Our work extends these insights to video generation, revealing that queries play a more complex role than previously understood - affecting both motion and identity. Additionally, we find that capturing motion patterns in videos requires more denoising steps compared to image generation, highlighting fundamental differences between these domains.
Extended Attention Sharing. When using text-to-image models to generate (Wu et al., 2023; Ceylan et al., 2023; Khachatryan et al., 2023) or modify a video (Geyer et al., 2023), an extended self-attention block (Wu et al., 2023) is often employed to share keys and values across different frames, enabling them to draw visual appearances from each other and enhance consistency. Beyond cross-frame consistency, it has been used to inject consistent identities from a source image to video (Xu et al., 2023; Hu et al., 2023; Chang et al., 2023; Tu et al., 2023), maintain appearance in layout editing (Cao et al., 2023; Avrahami et al., 2024), combine appearances (Alaluf et al., 2023), for personalization (Gal et al., 2024; Zeng et al., 2024) and style transfer (Hertz et al., 2023).
3 Analysis: The role of Query features in text-to-video generation
To study the role of Q-features in text-to-video (T2V) models, we design an injection experiment. The idea is simple: we take a source video $V_{S}$ with a known text description $\tau_{S}$, and also generate a target video $V_{T}$, using a text prompt $\tau_{T}$. We then record Q features from $V_{S}$, inject them into the generation of $V_{T}$, and analyze the effect of that injection.
More specifically, Given the video $V_{S}$ we add noise to it with a level that correspond to some noisy step $t$, yielding a noisy latent $z_{t}$. We then perform a single DDPM denoising step (with a $50$ step schedule), and record the spatial Q features of all the self attention layers of the diffusion model. We repeat this $20$ times for various noise levels. At the end of this procedure, we have a sequence of $20$ $Q$ tensors ($Q_{S}(50),Q_{S}(49),...,Q_{S}(30)$ DDPM steps corresponding to t=1000, 980, 960, …, 600). Finally, we generate a new video $V_{T}$ with its prompt $\tau_{T}$, while injecting $Q_{S}(t)$ tensors at the first k DDPM denoising steps. We vary the amount of denoising steps that receive Q-injection: From no injection at all (0%) to injecting $Q$ through 40% of DDPM steps.
To understand the effect of Q-injection, we now measure similarity between source and target videos in two aspects: Identity Leakage and Motion Fidelity. Identity Leakage measures the mean DINO similarity between the frames of the source $V_{S}$ and target $V_{T}$ videos. Motion Fidelity Yatim et al. (2024) measures cross-correlation between point tracks in source and target videos:
$\frac{1}{m}\sum_{\tilde{p}\in\tilde{\mathcal{P}}}\underset{p\in\mathcal{P}}{\text{max}}\ \textbf{corr}(p,\tilde{p})+\frac{1}{n}\sum_{{p}\in{\mathcal{P}}}\underset{\tilde{p}\in\tilde{\mathcal{P}}}{\text{max}}\ \textbf{corr}(p,\tilde{p})$ (1)
where $\mathcal{P}=\{p_{1},\dots,p_{n}\},\tilde{\mathcal{P}}=\{\tilde{p}_{1},\dots,\tilde{p}_{m}\}$ are point tracks in source and target videos.
We present our analysis results in Fig. 2. We analyze two Q-injection setups: one where the source video prompt $\tau_{S}$ and target prompt $\tau_{T}$ share the same subject (purple) and another where they have different subjects (green). We discuss three key findings, highlighting the differences in how Q vectors behave in T2V models compared to T2I models.
First, Fig. 2 (bottom-left), illustrates that Motion Fidelity grows monotonically with Q injection duration and reaches high similarity at 40% injection. We find that, unlike images–where structure is established early in the denoising process–videos require significantly more steps to set the motion structure. Second, more surprising is the top-left panel. It reveals that identity similarity also grows monotonically with duration of injecting Q. This suggests that video generation models also encode identity information into the Q vectors–an intriguing shift from its traditionally assumed role in T2I models. Third, we observe an interesting phenomenon: when $\tau_{S}$ and $\tau_{T}$ use the same subject, the target video often features a subject with an appearance identical to that of the source video while maintaining the background specified in $\tau_{T}$. This identity “leakage” is significantly less pronounced when $\tau_{T}$ and $\tau_{S}$ feature different subjects. Qualitatively, in Fig. 2 (right), we present a source video of “A horse galloping in the savanna” with a distinct white spot on its back. Notice that as the number of injection steps increases, the target horse (purple) becomes more similar to the source horse while preserving the cloud background specified in the prompt. On the other hand, when $\tau_{T}$ describes a giraffe, its shape and motion become more similar to those of the source horse, but its appearance remains that of a giraffe. This suggests a clear distinction in Q-injection behavior: it transfers appearance and motion when the subjects match, while for different subjects, the transfer primarily affects motion characteristics.
@@ 3 Analysis: The role of Query features in text-to-video generation
Finally, we examine the effect of Q-injection when the target consists of multiple videos sharing generation features. Feature sharing within a batch is a widely used approach in T2I generation, aiding tasks such as consistent character generation  (Tewel et al., 2024) and editing  (Cao et al., 2023). A common technique is the extended attention mechanism, which modifies the model’s self-attention layers to attend across elements of the multiple elements in the same batch. Here, we explore the impact of Q-injection when generating a batch of 3 videos at a time and utilizing the shared Extended Attention between the video shots. In Fig. 3, we include the Extended Attention curve to visualize motion fidelity in this scenario. Our results show that extended attention induces a ”motion freeze” effect, requiring Q to be injected for more steps to achieve the same motion as the source video. Moreover, since Q-injection in the Extended Attention setting requires more steps, it also introduces identity information, making it difficult to disentangle motion injection from identity injection.

### v3

@@ Abstract
Text-to-video diffusion models have shown remarkable progress in generating coherent video clips from textual descriptions. However, the interplay between motion, structure, and identity representations in these models remains under-explored. Here, we investigate how self-attention query ($Q$) features simultaneously govern motion, structure, and identity and examine the challenges arising when these representations interact. Our analysis reveals that Q affects not only layout, but that during denoising Q also has a strong effect on subject identity, making it hard to transfer motion without the side-effect of transferring identity. Understanding this dual role enabled us to control query feature injection (Q injection) and demonstrate two applications: (1) a zero-shot motion transfer method – implemented with VideoCrafter2 and WAN 2.1 – that is 10$\times$ more efficient than existing approaches, and (2) a training-free technique for consistent multi-shot video generation, where characters maintain identity across multiple video shots while Q injection enhances motion fidelity.
1 Introduction
Video generation from text is at the forefront of generative AI. Despite progress in controlling entities in video, several major challenges remain, including generating natural, engaging motion and preserving consistent identity of entities throughout the video. These two goals often form a trade-off: It is easy to preserve consistency if motion is strongly limited, and making entities move makes it harder to enforce consistency, as the appearance of an entity changes. A major challenge lies in understanding how motion and identity are represented in various components of video generation models and how to effectively control them.
This limited understanding hinders downstream applications in video generation. For instance, while many current motion transfer approaches Yatim et al. (2024); Zhao et al. (2024); Wang et al. (2024a); Liu & Vahdat (2025) rely on tuning or test-time optimization, there is growing interest in developing inference-time methods, similar to how text-to-image layout transfer already operates through feature manipulation Cao et al. (2023); Alaluf et al. (2023). Better model understanding could lead to further progress in this direction. As another example, consider consistent characters in multi-shot video generation, where the goal is to preserve consistency of character identity and appearance across shots. Image-based models tackle this through feature-sharing, but applying the same ideas to video leads to loss of motion because the shared features encode both identity and motion.
@@ 1 Introduction
To understand representation of motion and identity in a video model, let us first draw analogies from text-to-image (T2I) models which are better understood. Motion can be viewed as a ”3D-shape” in the 3D tensor defined by a sequence of frames, so it is natural to look at representation of shape and structure in image generation models. Previous works showed that diffusion-based T2I models establish the layout of the generated image in early steps (Patashnik et al., 2023). Several studies Cao et al. (2023); Alaluf et al. (2023); Tewel et al. (2024) analyzed how different model components determine this structure. They showed that self-attention queries (Q) encode structural layout information, and injecting queries between images during generation “copies” shape while preserving appearance. As motion is the video equivalent of structure, we investigate the relationship between query vectors, motion, and identity in video generation.
@@
For motion transfer, we find that injecting Q features from a source video during denoising of a generated video allows transferring motion to a new video in a zero-shot manner (without extra fine-tuning or optimization). Our simple pipeline achieves generation quality close to leading methods, while being 10$\times$ more efficient than existing approaches.
@@
Our main contributions: 1) We provide a systematic analysis of Q-features in text-to-video diffusion models, revealing their dual role in encoding both motion and identity information, with effects persisting longer into the denoising process. 2) We introduce “Motion by Queries”, an efficient approach that enables zero-shot motion transfer in both UNeT and Diffusion-Transformer architectures. 3) We present a training-free method for consistent multi-shot video generation that balances character consistency and motion quality.
@@ 2 Related work
Motion Transfer in Video Generation: Motion transfer guides video generation using source video motion patterns, spanning camera movements and object deformations while maintaining structural integrity. Common approaches involve model adaptation (fine-tuning temporal attention (Jeong et al., 2024; Ren et al., 2024), LoRAs (Zhao et al., 2024), motion embeddings (Wang et al., 2024a)) or test-time optimization (Yatim et al., 2024). Concurrent noise-warped methods Liu & Vahdat (2025); Burgert et al. (2025) require weeks of training on 8/64 GPUs to inject motion via warped noise, and are out of the scope of our zero-shot setup.  In contrast, our method enables zero-shot motion transfer via query injection without extra training or optimization. In Ling et al. (2025), a zero-shot approach is also employed using temporal attention maps, but incurs $\times$12 denoising overhead. Our work analyzes motion in a broader context through spatial Q features, providing principles applicable across video generation tasks and models.
Figure 2: Same-class motion transfer suffers from identity leakage (purple), worsening with increased Q injection. Cross-class transfer (green) achieves reasonable separation at 40% injection, where motion quality is also preserved. The leftmost purple data point shows results for random same-class images (no injection). Quantitative results use motion-transfer data from Wang et al. (2024a). For Illustration we use frames from the videos in Fig. 1.
Video Editing: Video editing modifies specific video attributes while preserving others, rather than replicating motion patterns. Bai et al. (2024) also use feature injection for edit and do not address our broader Q injection context. Zhao et al. (2023) combine ControlNet and CLIP. Image-based methods (Qi et al., 2023; Liu et al., 2024; Shin et al., 2024; Cong et al., 2023; Wu et al., 2023) struggle with coherence, as they lack a temporal model. Structure and Appearance in Text-to-Image Models: Early work in text-to-image showed that Keys and Queries guide structure (“Where”), while Values set appearance (“What”) Tewel et al. (2023). Further studies established that the layout is set early in denoising (Patashnik et al., 2023), and self-attention queries encode structural layout  (Alaluf et al., 2023; Tewel et al., 2024). We extend these findings to video, showing queries play a more complex role than previously understood - affecting both motion and identity. Additionally, we find that capturing motion patterns in videos requires more denoising steps compared to images, highlighting fundamental differences between these domains. Extended Attention Sharing: When using text-to-image models to generate (Wu et al., 2023; Ceylan et al., 2023; Khachatryan et al., 2023) or modify a video (Geyer et al., 2023), an extended self-attention block (Wu et al., 2023) is often employed to share keys and values across different frames, enabling them to draw visual appearances from each other and enhance consistency. Beyond cross-frame consistency, it has been used to inject consistent identities from a source image to video (Xu et al., 2023; Hu et al., 2023; Chang et al., 2023; Tu et al., 2023), maintain appearance in layout editing (Cao et al., 2023; Avrahami et al., 2024), combine appearances (Alaluf et al., 2023), for personalization (Gal et al., 2024; Zeng et al., 2024) and style transfer (Hertz et al., 2023). Consistent generation aims to maintain consistent subjects across outputs from a generative model, and has mainly studied in text-to-image tasks. A common approach is to use personalization (Gal et al., 2022; Ruiz et al., 2022) to promote consistency, either through inpainting with a personalized model (Jeong et al., 2023), personalized LoRA models (Ryu, 2023), or training LoRAs for large, semi-consistent clusters (Avrahami et al., 2023). Alternatively, encoders can inject identity at inference (Ye et al., 2023; Wei et al., 2023; Gal et al., 2023), but require large pre-training and struggle with new domains. Storyboard-based methods (Feng et al., 2023; Liu et al., 2023), and video personalization methods Jiang et al. (2024); Yuan et al. (2025); Wu et al. (2024); Zhang et al. (2025) face similar limits. Recent work  (Tewel et al., 2024; Fan et al., 2024) achieves character consistency without personalization or tuning, using feature sharing across images.
@@
3 Analysis: Query features in text-to-video generation
To study the role of Q-features in text-to-video (T2V) models, we design an injection experiment. The idea is simple: we take a source video $V_{S}$ with a known text description $\tau_{S}$, and generate a target video $V_{T}$ using a prompt $\tau_{T}$. We then record Q features from $V_{S}$, inject them into the generation of $V_{T}$, and analyze the effect of that injection. We base our main analysis on VideoCrafter2 Chen et al. (2024) model, and later test Q-injection in other architectures for generality.
More specifically, given the video $V_{S}$, we add noise at a level corresponding to a noisy step $t$, yielding a noisy latent $z_{t}$. We then perform a single DDPM denoising step (with a $50$-step schedule) and record the Q features from all spatial self-attention layers of the diffusion model. This is repeated $20$ times for various noise levels, resulting in a sequence of $20$ $Q$ tensors ($Q_{S}(50),Q_{S}(49),...,Q_{S}(30)$, corresponding to DDPM steps $t=1000,980,...,600$). Finally, we generate a new video $V_{T}$ with prompt $\tau_{T}$, while injecting $Q_{S}(t)$ tensors at the first $k$ DDPM denoising steps. We vary the amount of steps receiving Q-injection: From none (0%) up to 40% of DDPM steps.
To understand the effect of Q-injection, we measure similarity between source and target videos in two aspects: Identity Leakage and Motion Fidelity. Identity Leakage measures mean DINO similarity between the frames of source $V_{S}$ and target $V_{T}$ videos. Motion Fidelity Yatim et al. (2024) measures cross-correlation between point tracks in source and target videos:
$\frac{1}{m}\sum_{\tilde{p}\in\tilde{\mathcal{P}}}\underset{p\in\mathcal{P}}{\text{max}}\ \textbf{corr}(p,\tilde{p})+\frac{1}{n}\sum_{{p}\in{\mathcal{P}}}\underset{\tilde{p}\in\tilde{\mathcal{P}}}{\text{max}}\ \textbf{corr}(p,\tilde{p})$
where $\mathcal{P}=\{p_{1},\dots,p_{n}\},\tilde{\mathcal{P}}=\{\tilde{p}_{1},\dots,\tilde{p}_{m}\}$ are point tracks in source and target videos. Error bars: are standard-error-of-the-mean (S.E.M) to show the significance of our findings.
Fig. 2 shows our results. We consider two Q-injection setups: one where the source and target prompts share the same subject (purple), and one where they differ (green). We highlight three key differences in how Q injection behaves in T2V models versus T2I models.
Figure 4: (click-to-view-online) Qualitative Results, Motion Transfer (VideoCrafter2). Each pair of rows are frames from source (pair-top) real video and target (pair-bottom) generated video. Transferring Motion by Queries allows to use source videos to inject camera motion (top), non-rigid movement (middle), and combinations of movements (bottom). See the supplemental for videos and more examples.
First, Fig. 2 (bottom-left), shows that Motion Fidelity increases with Q-injection duration, reaching high similarity at 40% injection. We find that, unlike images–where structure is established early in the denoising process–videos require significantly more steps to set the motion structure. Second, more surprising is the top-left panel. It reveals that identity similarity also increases with the duration of injecting Q. This suggests that video generation models also encode identity information into the Q vectors–an intriguing shift from its traditionally assumed role in T2I models. Third, we observe an interesting phenomenon: when $\tau_{S}$ and $\tau_{T}$ use the same subject, the target video often features a subject with an appearance identical to that of the source video while maintaining the background specified in $\tau_{T}$. This identity “leakage” is significantly less pronounced when $\tau_{T}$ and $\tau_{S}$ feature different subjects. Qualitatively, in Fig. 2 (right), we present a source video of “A horse galloping in the savanna” with a distinct white spot on its back. Notice that as the number of injection steps increases, the target horse (purple) becomes more similar to the source horse while preserving the cloud background specified in the prompt. On the other hand, when $\tau_{T}$ describes a giraffe, its shape and motion become more similar to those of the source horse, but its appearance remains a giraffe. This suggests a clear distinction in Q-injection behavior: it transfers appearance and motion when the subjects match, while for different subjects, the transfer primarily affects motion characteristics.
@@ 3 Analysis: Query features in text-to-video generation
Finally, we examine the effect of Q-injection when the target consists of multiple videos sharing generation features. Feature sharing within a batch is widely used in T2I generation, aiding tasks like consistent character generation (Tewel et al., 2024) and editing (Cao et al., 2023). A common technique is the extended attention mechanism, which modifies self-attention layers to attend across elements of multiple videos in the same batch. Here, we explore the impact of Q-injection when generating a batch of 3 videos at a time and using shared Extended Attention between shots. In Fig. 3, we include the Extended Attention curve to visualize motion fidelity in this scenario. Our results show that extended attention induces a ”motion freeze” effect, requiring Q to be injected for more steps to match the source video’s motion. Moreover, since Q-injection in this setting needs more steps, it also introduces identity information, making it difficult to disentangle motion from identity injection.
Generality Across Architectures. We extended our core Q-injection experiment of Fig. 1 to several T2V models: (1) WAN 2.1 Wang et al. (2025), a SoTA DiT model; (2) T2V-Turbo-V2 Li et al. (2024a), a fast sampling model, using as few as 8 denoising steps; and (3) LTX-Video HaCohen et al. (2024), a fast DiT model. As in our main findings, Q-injection led to identity leakage for same-subject prompts and primarily transferred motion for different-subject prompts. We also verified that without Q-injection, source and target prompts produced visually distinct outputs. Qualitative results are provided in the supplemental and Fig. 8. See Appendix for implementation details.